OpenStudy (anonymous):

find the radius of a circle with circumferene of 50cm

OpenStudy (shinalcantara):

remember that circumference is \[C = 2 \pi r\] since you've got the circumference as 50cm, then you can find r by substituting the value to the equation and you'll have r as \[r = \frac{ C }{ 2 \pi }\]

OpenStudy (igreen):

^ Just follow what @shinalcantara gave you. \(r = \dfrac{C}{2 \pi}\) Plug in the circumference and pi: \(r = \dfrac{50}{2(3.14)}\) Multiply: \(r = \dfrac{50}{6.28}\) Can you divide that? @ogolotswe

OpenStudy (anonymous):

yes thank u

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @ogolotswe